If the equation x^2 – x – p =0 and x^2 - 2xp - 12  =0 have a common root ,find it

Grade:11

1 Answers

Y RAJYALAKSHMI
9 years ago
x^2-x-p=0; p = (-x^2 +x)
substitute p in 2nd equation
x^2-2x(-x^2+x)-12=0
2x^3-x^2-12=0
solving this eqn we get x = 2 and other roots are unreal
so, the common root is 2